# Learning in Focus: Detecting Behavioral and Collaborative Engagement Using Vision Transformers

### v1

@@ Abstract
In early childhood education, accurately detecting behavioral and collaborative engagement is essential for fostering meaningful learning experiences. This paper presents an AI-driven approach that leverages Vision Transformers (ViTs) to automatically classify children’s engagement using visual cues such as gaze direction, interaction, and peer collaboration. Utilizing the Child-Play gaze dataset, our method is trained on annotated video segments to classify behavioral and collaborative engagement states (e.g., engaged, not engaged, collaborative, not collaborative). We evaluated three state-of-the-art transformer models: Vision Transformer (ViT), Data-efficient Image Transformer (DeiT), and Swin Transformer. Among these, the Swin Transformer achieved the highest classification performance with an accuracy of 97.58%, demonstrating its effectiveness in modeling local and global attention. Our results highlight the potential of transformer-based architectures for scalable, automated engagement analysis in real world educational settings.
Index Terms:
Behavioral engagement, collaborative engagement, vision transformers (ViTs), Swin, DeiT.
@@ I Introduction
In this paper, we introduce “Learning in Focus”, a framework for identifying behavioral and collaborative involvement in children using video. We define behavioral engagement as a child’s focused attention and on-task conduct during learning, and collaborative engagement as the level of interactive participation and joint problem solving within a group (see Fig. 1. To illustrate model outputs while maintaining privacy and avoiding the exposure of identifiable features, we displayed only Canny edge-detected versions of the gaze frames in our figures. The original annotated images themselves were never shown or directly revealed). Using the ChildPlay gaze dataset [22], we process our model using a Vision Transformer that has been trained on large visual datasets. The primary aim of the project is to
@@
  • Evaluate and compare the effectiveness of several transformer-based architectures, such as ViT, Swin, and DeiT, in classifying engagement and collaboration.
@@
The rest of the paper is organized as follows. Section II presents related work on vision-based engagement detection and transformer models used in classroom settings. Section III describes the methodology, including the Swin Transformer architecture and its hierarchical components. Section IV discusses experimental setup, dataset preparation, evaluation metrics and compares the performance of Swin, ViT, and DeiT models. Finally, Section V concludes the paper and outlines future directions for research in automated engagement recognition using visual data.
@@ II Related works
Additionally, educational psychology defines engagement as a multidimensional construct[19], with behavioral engagement being a strong predictor of academic success[20]. Collaborative engagement emphasizes shared focus, coordination, and responsiveness among peers[13]. Recent research aligns these psychological definitions with computer vision features such as gaze, posture, and interaction sequences[18], shifting the measurement paradigm from self-reports to behaviorally grounded annotation.
Despite advancement, obstacles persist. Many data sets lack fine-grained frame-level labels [15], which limits model generalization. Existing research often focuses on small groups, but real classrooms confront occlusion and scalability challenges[21]. Deep models struggle with delicate behaviors and class imbalance, and few work addresses interpretability [21]. Multimodal fusion is promising, but it is computationally demanding, which limits its practical application. Although vision and deep learning have improved engagement detection, important gaps remain in dataset quality, scalability, and model interpretability. Our approach solves these issues using ViTs to provide a scalable and explainable framework for detecting collaborative and behavioral engagement in children via gaze data.
@@ III-A Background
TABLE I: Detailed Stage-wise Configuration and Description of Swin-B (Base) Transformer [27].
@@
1 $56\times 56$ Concat $4\times 4$, 128-d, LayerNorm (LN) $7\times 7$ 128 4 2 Image is split into $4\times 4$ patches and are embedded; local attention is applied to capture the fine-grained visual cues.
2 $28\times 28$ Concat $2\times 2$, 256-d, LayerNorm (LN) $7\times 7$ 256 8 2 Spatial resolution is halved while increasing channel depth; captures broader regional features using window-based attention.
3 $14\times 14$ Concat $2\times 2$, 512-d, LayerNorm (LN) $7\times 7$ 512 16 18 Feature map is compressed further; 18 deep transformer blocks allow modeling of high-level context and interactions.
4 $7\times 7$ Concat $2\times 2$, 1024-d, LayerNorm (LN) $7\times 7$ 1024 32 2 Final global representation is formed; deep embeddings summarize the entire visual context for classification.
@@ IV Experimental Results and Discussion
This section presents and analyzes the experimental results obtained from training and evaluating the ViT, DeiT, and Swin Transformer models, focusing on their performance across various metrics, learning behaviors, and classification tasks.
@@ IV-A2 Metrics
Not Collaborative (engaged but not socially) The child is engaged with the activity but not interacting with the partner. They show no shared attention – interest is only in the object itself, not in cooperating. Even though they’re focused (engaged), they are not collaborating. Here, the social partner is available but the child remains in solo mode. [Uncaptioned image][Uncaptioned image]
@@ IV-A3 Hyperparameters
To enable an accurate comparison, all three models (ViT, Swin, and DeiT) were trained under the same parameters. As can be seen from Table I, all models employed the same learning rate of 2e-4, batch size of 32, and weight decay of 1e-2. Each model was trained for 20 epochs with the AdamW optimizer, which is noted for its stability and effectiveness when training transformer architectures. To improve model generalization and address class imbalance, a shared set of data augmentation strategies was used. These included random resizing cropping, horizontal flipping, random rotation, color jittering, tensor conversion, and normalizing. By using similar hyperparameters and augmentation techniques, the evaluation focuses solely on architectural differences, guaranteeing that any performance variances detected are due to model capabilities rather than tuning inconsistencies.
@@
Model Precision Recall F1 Score Accuracy Parameters
@@ IV-B Results
We conducted experiments to compare the performance of three cutting-edge transformer-based models: Vision Transformer (ViT) [23], Shifted Window Transformer (Swin) [27] and Data-efficient Image Transformer (DeiT) [28] using a labeled dataset consisting of 4,537 tagged images. All three models were fine tuned with pretrained weights derived from large scale image datasets. It allows us to use rich feature representations while achieving robust performance on a relatively small dataset.
@@
(c) Swin
Figure 3: Learning curves of transformer models.
@@
(c) Swin
Figure 4: Confusion matrices of different transformer models.
(a) ViT
(b) DeiT
(c) Swin
@@
IV-B1 Learning Curve
Fig 3 (a), (b), (c) illustrate the learning curves of the transformer models, which give information about the training dynamics and generalization behavior over 20 epochs. All three models show a significant decreased trend in both training and validation loss, demonstrating good learning during optimization. In the case of ViT and DeiT, the validation loss closely follows the training loss, with rare fluctuations, indicating steady training with minor deviations most likely due to data complexity. The Swin Transformer, shown in Fig 3(c), has a more constant gap between training and validation loss while keeping a lower overall validation loss, indicating higher generalization and less overfitting. Importantly, no model shows signs of divergence or severe overfitting during the training period. These graphs show that all three models were effectively trained, with Swin exhibiting slightly more robust and consistent convergence behavior than ViT and DeiT.
IV-B2 Statistical Comparison
Table III presents a comparison of three transformer-based models, Vision Transformer (ViT), Data-efficient Image Transformer (DeiT), and Swin Transformer (Swin) evaluated on an classifying engagement detection task. Among the three, the Swin Transformer outperformed the others, achieving the highest accuracy of 97.58%, along with strong precision (0.96), recall (0.98) and F1 scores (0.97). This performance indicates the effectiveness of its hierarchical attention mechanism and localized windowing strategy, which helped to capture spatial patterns more efficiently in the dataset.
Following closely, DeiT achieved competitive performance with an accuracy of 97.47% and the precision of 0.9623. This shows that its training efficiency and knowledge distillation technique improved its accuracy and generalization. ViT also performed well, with an accuracy of 97.25%, demonstrating the effectiveness of its patch-based self-attention mechanism in handling visual data. While ViT and DeiT both have around 86 million parameters, Swin has slightly more at 88 million due to its more complex architecture.
Overall, all three transformer models showed high performance for image classification tasks. The subtle differences in their performance highlight the strengths of each architecture. However, the Swin Transformer exhibited a slight but consistent advantage in overall accuracy, making it the top performing model in this comparison.
TABLE IV: Qualitative Comparison of Various Methods Across Some Key Elements.
Features Chen et al. (MDNN) [2] Rogat et al. (2022 Rubric) [3] Tsai et al. (2020 OpenPose) [7] Abdelkawy et al. [10] Alarefah et al. [8] Wu et al. (CMOSE) [11] Ours (Learning in Focus)
Vision-Based
Uses Gaze
Collaborative Engagement Detection
Behavioral Engagement Detection
Both Collaborative & Behavioral
Multiclass Labels
Temporal Tracking
IV-B3 Confusion Matrix
@@
Fig 4 (a), (b), (c) show the confusion matrices for ViT, DeiT, and Swin Transformers, respectively, highlighting each model’s ability to classify engagement and collaboration levels. All three models successfully identified the majority of samples across classes, with the Swin Transformer producing the most balanced and accurate results. ViT performed well overall, but there was some confusion between the “engaged” and “collaborative” classes. DeiT showed better precision, particularly in the “collaborative” category, despite minor misclassifications across all classes. Overall, while ViT and DeiT performed well, the Swin Transformer stood out due to its clearer separation and higher accuracy across all behavioral categories.
IV-B4 ROC Curve
Fig 5 (a), (b), (c) depicts the ROC curves for the ViT, DeiT, and Swin transformer models, which provide classification performance across four behavior categories: engaged, not_engaged, collaborative, and not_collaborative. All three models had an AUC value of 1.00 for each class and the micro-average, showing high discriminative ability and a good combination of sensitivity and specificity. The curves are tightly grouped in the upper left corner, indicating a high true positive rate with few false positives. Among the three, the Swin Transformer had somewhat more stable and smoother curves, indicating more confident and consistent predictions across all classes. These findings support the robustness and reliability of all three transformer models, with Swin demonstrating a minor performance advantage in terms of ROC behavior.
IV-B5 Qualitative Comparison
Table IV compares engagement detection systems based on essential elements like vision-based input, gaze utilization, collaborative and behavioral detection, multiclass classification, and temporal tracking. Most methods rely on vision-based tactics, with only a few incorporating gaze or supporting both modes of involvement. Behavioral detection is more widespread, but collaborative participation receives less attention. Our approach (Learning in Focus) is the only one that addresses all essential areas except temporal modeling, demonstrating its greater breadth than previous studies.
@@ V Conclusion
In this paper, we explored how transformer based models can be used to detect behavioral and collaborative engagement in learning environments. Using cropped images extracted from the ChildPlay Gaze dataset, we trained and evaluated models like ViT, DeiT, and Swin Transformer to recognize different types of engagement. Among them, the Swin Transformer stood out for its ability to capture both fine-grained and broad visual patterns, leading to the best overall performance. In the next phase, we plan to advance from static image classification to a more dynamic video-based understanding. By incorporating transformer-based models like the Video Swin Transformer, we aim to effectively capture temporal patterns such as gaze shifts and interaction cues over time. This will enable behavior classification at the frame level and support real-time analysis as well. Overall, this pipeline will help us build a context-aware and scalable system that better aligns with real-world scenarios and improves the automated detection of engagement and collaboration.

### v2

@@ Abstract
In early childhood education, accurately detecting collaborative and behavioral engagement is essential to foster meaningful learning experiences. This paper presents an AI-driven approach that leverages Vision Transformers (ViTs) to automatically classify children’s engagement using visual cues such as gaze direction, interaction, and peer collaboration. Utilizing the Child-Play gaze dataset, our method is trained on annotated video segments to classify behavioral and collaborative engagement states (e.g., engaged, not engaged, collaborative, not collaborative). We evaluated six state-of-the-art transformer models: Vision Transformer (ViT), Data-efficient Image Transformer (DeiT), Swin Transformer, VitGaze, APVit and GazeTR. Among these, the Swin Transformer achieved the highest classification performance with an accuracy of 97.58%, demonstrating its effectiveness in modeling local and global attention. Our results highlight the potential of transformer-based architectures for scalable, automated engagement analysis in real world educational settings.
@@ I Introduction
To overcome these challenges, we introduced “Learning in Focus”, a framework for identifying behavioral and collaborative involvement in children using video. We define behavioral engagement as a child’s focused attention and on-task conduct during learning, and collaborative engagement as the level of interactive participation and joint problem solving within a group (see Fig. 1. To illustrate model outputs while maintaining privacy and avoiding the exposure of identifiable features, we displayed only Canny edge-detected versions of the gaze frames in our figures. The original annotated images themselves were never shown or directly revealed). Using the ChildPlay gaze dataset [22], we process our model using a Vision Transformer that has been trained on large visual datasets. The primary aim of the project is to
@@
  • Evaluate and compare the effectiveness of several transformer-based architectures, such as ViT, Swin, and DeiT, VitGaze, APVit and GazeTR in classifying engagement and collaboration.
@@
The rest of the paper is organized as follows. Section II presents related work on vision-based engagement detection and transformer models used in classroom settings. Section III describes the methodology, including the Swin Transformer architecture and its hierarchical components. Section IV discusses experimental setup, dataset preparation, evaluation metrics and compares the performance of Swin, ViT, DeiT, VitGaze, APVit and GazeTR models. Finally, Section V concludes the paper and outlines future directions for research in automated engagement recognition using visual data.
@@ II Related works
Additionally, educational psychology defines engagement as a multidimensional construct[19], with behavioral engagement being a strong predictor of academic success[20]. Collaborative engagement emphasizes shared focus, coordination, and responsiveness among peers[13]. Recent research aligns these psychological definitions with computer vision features such as gaze, posture, and interaction sequences[18], shifting the measurement paradigm from self-reports to behaviorally grounded annotation. Despite advancement, obstacles persist. Many data sets lack fine-grained frame-level labels [15], which limits model generalization. Existing research often focuses on small groups, but real classrooms confront occlusion and scalability challenges[21]. Deep models struggle with delicate behaviors and class imbalance, and few work addresses interpretability [21]. Multimodal fusion is promising, but it is computationally demanding, which limits its practical application. Although vision and deep learning have improved engagement detection, important gaps remain in dataset quality, scalability, and model interpretability. Our approach solves these issues using ViTs to provide a scalable and explainable framework for detecting collaborative and behavioral engagement in children via gaze data.
@@ III-A Background
TABLE I: Stage-wise Configuration and Description of Swin-B (Base) Transformer [27].
@@
1 $56\times 56$ Concat $4\times 4$, 128-d, LN $7\times 7$ 128 4 2 Splits image into patches and applies local attention.
2 $28\times 28$ Concat $2\times 2$, 256-d, LN $7\times 7$ 256 8 2 Reduces resolution and learns broader regional features.
3 $14\times 14$ Concat $2\times 2$, 512-d, LN $7\times 7$ 512 16 18 Deep blocks capture high-level global interactions.
4 $7\times 7$ Concat $2\times 2$, 1024-d, LN $7\times 7$ 1024 32 2 Forms final feature map for global representation.
@@ IV Experimental Results and Discussion
This section presents and analyzes the experimental results obtained from training and evaluating the ViT [23], DeiT [28], VitGaze [29], APVit [30] and GazeTR [31] and Swin [27] Transformer models, focusing on their performance across various metrics, learning behaviors, and classification tasks.
@@ IV-A2 Metrics
Not Collaborative (engaged but not socially) The child is engaged with the activity but not interacting with the partner. They show no shared attention. Even though they are focused (engaged), they are not collaborating. [Uncaptioned image][Uncaptioned image]
@@ IV-A3 Hyperparameters
To enable an accurate comparison, all six models (ViT, Swin, VitGaze, APVit and GazeTR and DeiT) were trained under the same parameters. As can be seen from Table I, all models employed the same learning rate of 2e-4, batch size of 32, and weight decay of 1e-2. Each model was trained with the AdamW optimizer, which is noted for its stability and effectiveness when training transformer architectures. To improve model generalization and address class imbalance, a shared set of data augmentation strategies was used. These included random resizing cropping, horizontal flipping, random rotation, color jittering, tensor conversion, and normalizing. By using similar hyperparameters and augmentation techniques, the evaluation focuses solely on architectural differences, guaranteeing that any performance variances detected are due to model capabilities rather than tuning inconsistencies.
@@
Model Precision Recall F1 Score Accuracy Params
@@ IV-B Results
We conducted experiments to compare the performance of six cutting-edge transformer-based models: ViT [23], Swin [27], DeiT [28], VitGaze [29], APVit [30] and GazeTR [31] using a labeled dataset consisting of 4,537 tagged images. All the models were fine tuned with pretrained weights derived from large scale image datasets. It allows us to use rich feature representations while achieving robust performance on a relatively small dataset.
@@
(c) ViTGaze
(d) APvit
(e) GazeTR
(f) Swin
Figure 3: Confusion matrices of different transformer models where rows correspond to actual labels and columns to predicted labels.
@@
(c) ViTGaze
(d) APVit
(e) GazeTR
(f) Swin
@@
IV-B1 Statistical Comparison
Table III presents a comparative evaluation of six transformer-based architectures: ViT, DeiT, VitGaze, ApViT, GazeTR, and Swin across four key performance metrics. Among all models, the Swin Transformer achieves the highest overall performance, obtaining the top recall (0.9805), F1-score (0.9701), and accuracy (0.9758), demonstrating strong capability in capturing fine-grained engagement cues. DeiT records the highest precision (0.9623), highlighting its ability to minimize false positives while maintaining accuracy comparable to ViT. The standard ViT model also performs strongly, achieving an F1-score of 0.9583, indicating that baseline transformer encoders remain highly competitive.
ApViT shows balanced and reliable performance with an F1-score of 0.9482, positioning it between lightweight gaze-specific models and larger transformer variants. Gaze-focused models such as VitGaze and GazeTR also perform well. However, GazeTR demonstrates notably stronger performance, achieving a high recall of 0.9709 and an accuracy of 0.9570, which is substantially higher than the accuracy of VitGaze (0.8502). Overall, the results indicates that while several models perform competitively, hierarchical designs such as Swin offer the best overall balance of accuracy and robustness for holistic engagement recognition.
IV-B2 Confusion Matrix
@@
Fig 3 (a)–(f) show the confusion matrices for the ViT, DeiT, VitGaze, APVit, GazeTR and Swin Transformers highlighting each model’s ability to classify engagement and collaboration levels. All models successfully identified the majority of samples across classes, with the Swin Transformer producing the most balanced and accurate results. ViT performed well overall, but there was some confusion between the “engaged” and “collaborative” classes. DeiT showed better precision, particularly in the “collaborative” category, despite minor misclassifications across all classes.
Although ViTGaze benefits from a lightweight design, its confusion matrix shows comparatively higher misclassification, particularly between engaged, not engaged, and collaborative states. In contrast, APViT and GazeTR exhibit much cleaner class separability, with strong diagonal dominance, while Swin delivers the most balanced and consistent predictions overall.
IV-B3 ROC Curve
Fig 4 (a)–(f) display the ROC curves for the six transformer-based models across the four behavioral categories. Five of the models (ViT, DeiT, APViT, GazeTR, and Swin) achieve near-perfect discriminative performance, with AUC values approaching 1.00 across all classes. ViTGaze shows slightly lower AUC values (0.95–0.98), indicating reduced sensitivity and specificity compared to the others. Although all models produce clustered curves near the top-left region (indicating low false-positive rates and high true-positive rates), Swin and APViT exhibit the most stable and consistently smooth ROC profiles, reflecting stronger confidence across decision thresholds.
@@ V Conclusion
In this paper, we explored how transformer based models can be used to detect behavioral and collaborative engagement in learning environments. Using cropped images extracted from the ChildPlay Gaze dataset, we trained and evaluated models like ViT, DeiT, VitGaze, APVit, GazeTR and Swin Transformer to recognize different types of engagement. Among them, the Swin Transformer stood out for its ability to capture both fine-grained and broad visual patterns, leading to the best overall performance. In the next phase, we plan to advance from static image classification to a more dynamic video-based understanding. By incorporating transformer-based models like the Video Swin Transformer, we aim to effectively capture temporal patterns such as gaze shifts and interaction cues over time. This will enable behavior classification at the frame level and support real-time analysis as well. Overall, this pipeline will help us build a context-aware and scalable system that better aligns with real-world scenarios and improves the automated detection of engagement and collaboration.
Acknowledgement: This work was supported by the BRAINS Lab (Bioinspired Robotics, AI, Imaging & Neurocognitive Systems Laboratory) in the Department of Computer Science at The University of Alabama.